# Video-Based Detection and Analysis of Errors in Robotic Surgical Training

### v1

@@ IV Discussion
We developed and quantified an image-processing algorithm to detect collision errors in a ring tower transfer task. This allowed us to study how surgical residents improved over six months in terms of completion time and error metrics. We discovered that they improved their task completion time, and that, on average, this did not come at the expense of increasing the percentage of error time over the six months.
Our error detection algorithm achieved an accuracy and F1 score of approximately 95%, demonstrating its ability to automatically detect collision errors in the surgical training task. Several previous works investigated the possibility of such detection in other surgical tasks. The researchers in [39] automatically detected errors, such as multiple attempts and tool out of view, in suturing and needle passing tasks in the JIGSAWS dataset. They used neural networks to classify surgical gestures as correct or erroneous, achieving F1 scores ranging between 75%-95%. Shao et al. [37] detected errors in both these tasks, and reported results on the frame level, as in our work. They achieved F1 and accuracy values of approximately 70%, and presented very efficient computation, allowing for real-time analysis. While showing promising results, both [39, 37] used neural networks and required the data to be labeled. Our algorithm did not include any prior indications of when errors were made and can serve as a tool for labeling the data for future training of networks. The high accuracy and F1 scores show the reliability of such labeling. In addition to allowing us to study the residents’ errors, we will use this now labeled dataset for future analyses.
While there was no labeling of the errors themselves, the algorithm did use manual event segmentation, allowing us to analyze the interaction with each of the four towers separately. However, there are potential ways to overcome the lack of such segmentation. First, the entire task video could have been analyzed without separation based on the tower. Frames with tower movements would likely still be characterized by higher frequencies in the velocity magnitude signal. We hypothesize that the algorithm would still be successful, but anticipate that the performance might be impacted by noisier velocity signals. Second, it may be possible to achieve the event segmentation automatically using the PSMs’ kinematic data. The task board was static, such that the beginning and end of the interactions with each tower would be near set coordinates. Additionally, the interaction with each tower involved only one hand, whereas movements between the towers involved both hands. Lastly, the order of the four towers remained constant. Hence, the beginning of the interaction with a tower could potentially be identified by correlating the arrival of the relevant PSM to the start area of the tower with near-zero velocity in the other PSM. Similarly, the end of the interaction with a tower would involve arriving at the end area of the tower, correlated with the other PSM beginning to move. These suggestions add computation, and would likely affect the performance to an extent; however, they do offer possibilities for fully automating the detection using our algorithm.
@@
Examining residents’ completion time together with error metrics allows for assessing their overall improvement as there is often a trade-off between the two [24, 27, 29, 42]. We found that, on average, the residents spent a relatively constant percentage of time colliding with the tower over the six months, whereas they improved their task completion time. Comparing each resident’s individual performance in the first and last of the 18 visits showed that 12 improved in both metrics, reflecting learning and improvement over this long-term study. However, there were six residents who had some increase in error percentage, raising the question of whether they improved, or rather traded off improvement in one aspect of the task for another.
We additionally learned that the characteristics of the errors changed over time. Our observed decrease in completion time and the number of errors, together with the relatively constant average error percentage, indicates that many residents began with making a larger number of shorter errors, and progressed to making fewer errors, each of which spanned a larger portion of the tower. This might reflect that at the beginning some residents were more careful, and with time, put a larger emphasis on completion time while making less of an effort to minimize collisions. Another possibility is that their increase in task speed over time, and the time needed to initiate a correction to the errors, may not have allowed for stopping the error quickly enough to decrease the error percentage [45].
The residents were asked to complete the task as quickly as possible while minimizing errors, however, received no feedback on their performance. Feedback has been shown to affect different aspects of performance [46], and it is therefore possible that different feedback strategies would have led to different results. For example, future works can test if penalizing errors leads to their decrease, and study if this comes at the expense of the completion time as expected based on the speed-accuracy trade-off [24, 27, 29, 42]. It is logical that given the lack of explicit feedback, the residents in our experiment preferred to finish faster due to their busy and demanding schedules, rather than take longer and attempt to reduce their errors.
@@
We found that the residents’ completion times decreased both between shifts and between the timings relative to the shifts. That is, despite being after a 26-hour shift, the residents still improved. Studies have shown that sleep can improve motor skills [47, 48], and it is possible that an even greater improvement would have been observed in the absence of fatigue. Furthermore, residents sometimes improved between the last visit of one month and the first visit of the next month, despite a month going by without them interacting with a surgical robot. This may indicate offline learning [48, 49], but future studies are needed to determine if such offline learning indeed took place.
To conclude, we developed an algorithm for the automatic detection of collision errors in a surgical training task. Beyond enabling us to study residents’ learning of the task over six months, this framework provides a labeled dataset for further analysis.
ACKNOWLEDGMENT
We thank all the residents who participated in our experiment, Yoav Farawi for completing the manual event segmentation, and Tami Matus for the endless and relentless administrative help at all hours.

### v2

@@ IV Discussion
We studied how surgical residents’ performance in the ring tower transfer task evolved over a six month period and assessed the effect of a 26-hour shift on the learning process. To this end, we developed an image-processing algorithm to detect the collision errors in this surgical training task. Using the results, we computed error metrics, and we additionally computed task completion time to provide a more comprehensive assessment of performance. We discovered that the residents improved their task completion time both in the sessions surrounding a single shift and across the six-month period. Furthermore, on average, this did not come at the expense of increasing their percentage of errors over time.
@@
Many various metrics can be computed to assess different aspects of performance. In Sharon et al. [sharon2025datasetanalysislongtermskill], kinematic metrics indicative of surgical skill, such as path length and rate of orientation change [sharon2021rate] were computed. Analyzing the residents’ errors in this work complements the previously evaluated kinematic metrics. We examined the residents’ error metrics together with their completion time as there is often a trade-off between the two [chien2010accuracy, fitts1954information, reis2009noninvasive, listman2021long]. We found that, on average, the residents spent a relatively constant percentage of time colliding with the tower over the six months, whereas they improved their task completion time. Comparing each resident’s individual performance in the first and last of the 18 visits showed that 12 improved in both metrics, reflecting learning and improvement over this long-term study. However, six residents showed an increase in error percentage, raising the question of whether they improved, or rather traded off performance in one aspect of the task for another. Further indications of performance improvement were documented in [sharon2025datasetanalysislongtermskill], which reported improvements in kinematic metrics over the six-month period.
@@
We also considered the potential effect of the 26-hour shift on the residents’ learning process. We found that their completion times decreased both between shifts and between the timings relative to the shifts. That is, despite being after a 26-hour shift, the residents still improved. Studies have shown that sleep can improve motor skills [walker2002practice, al2013sleep], and it is possible that an even greater improvement would have been observed in the absence of fatigue. The reduction in completion time is consistent with the reported improvement in path length across the timings relative to the shifts, whereas the kinematic metric rate of orientation change did not reflect this improvement [sharon2025datasetanalysislongtermskill]. Furthermore, residents sometimes improved their completion time between the last visit of one month and the first visit of the following month, despite not interacting with a surgical robot during the interim. This may indicate offline learning [al2013sleep, dayan2011neuroplasticity], but future studies are needed to determine if such offline learning indeed took place.
We additionally learned that the characteristics of the errors changed over the six months. Our observed decrease in completion time and number of errors, together with the relatively constant average error percentage, indicates that many residents began with making a larger number of shorter errors, and progressed to making fewer errors, each of which spanned a larger portion of the tower. This might reflect that at the beginning some residents were more careful, and with time, put a larger emphasis on completion time while making less of an effort to minimize collisions. Another possibility is that their increase in task speed over time, and the time needed to initiate a correction to the errors, may not have allowed for stopping the error quickly enough to decrease the error percentage [briere2011automatic].
The residents were asked to complete the task as quickly as possible while minimizing errors; however, they received no feedback on their performance. Feedback has been shown to influence various aspects of performance [kluger1996effects, 11202637], and it is therefore possible that different feedback strategies would lead to different results. Among the various potential applications of error detection in surgical tasks [shao2024think, sirajudeen2024deep, li2022runtime, xu2024sedmamba], one is the development of surgical training protocols. The design of effective training protocols involves many components, one of which is the performance feedback provided to the surgeon. As one illustration, Yang et al. [yang2017effectiveness] demonstrated that presenting users with videos of surgical errors made by other novices, alongside a video of an expert correctly performing the surgical task, resulted in better learning compared to watching only the expert performance. This demonstrates the importance of understanding how surgeons acquire robotic surgical skills and highlights the potential value of error-related feedback in such learning. Future studies can test whether penalizing errors leads to their decrease and examine whether this comes at the expense of the completion time, as expected based on the speed-accuracy trade-off [chien2010accuracy, fitts1954information, reis2009noninvasive, listman2021long]. It is logical that, given the lack of explicit feedback, the residents in our experiment preferred to finish faster due to their busy and demanding schedules rather than take longer and attempt to reduce their errors. However, as errors in surgery can cause serious injuries [sarker2005errors, dankelman2005systems, bosma2011incidence, etchells2003patient], there can be cases in which prioritizing error reduction over completion time could be beneficial.
To reduce reliance on manual annotation in this study, we developed an algorithm for the automatic detection of collision errors in the ring tower transfer task. Although we manually checked the algorithm and corrected the automatic labels when necessary, this was more consistent and far less time-consuming than manually labeling the data. We found that our algorithm achieved accuracy and F1 scores of approximately 95%, demonstrating its ability to automatically detect collision errors in the surgical training task. Several previous studies have developed algorithms for automatically detecting surgical errors [shao2024think, sirajudeen2024deep, li2022runtime, xu2024sedmamba]. For example, the researchers in [li2022runtime] automatically detected errors such as multiple attempts and tool out of view in suturing and needle passing tasks in the JIGSAWS dataset. They used neural networks to classify surgical gestures as correct or erroneous, achieving F1 scores ranging between 75%-95%. Shao et al. [shao2024think] detected errors in the same tasks, achieving accuracy and F1 scores of approximately 70% on the frame level, and presented very efficient computation, allowing for real-time analysis. While showing promising results, both [li2022runtime, shao2024think] used neural networks and required the data to be labeled. Our algorithm did not rely on any prior indications of when errors occurred and can serve as a tool for labeling the data for future training of networks. The high accuracy and F1 scores show the reliability of such labeling. In addition to allowing us to study the residents’ errors, we will use this labeled dataset for future analyses.
It is important to note that although the errors themselves were not labeled, the algorithm did use manual event segmentation, which allowed us to analyze the interaction with each of the four towers separately. This segmentation was conducted as part of a previous analysis of the data and was utilized in this study. However, we suggest potential ways to overcome the lack of such segmentation to enable further automation of the algorithm. First, the entire task video could have been analyzed without separation based on the tower. Frames with tower movements would likely still be characterized by higher frequencies in the velocity-magnitude signal. We hypothesize that the algorithm would still be successful, but anticipate that the performance might be impacted by noisier signals. Second, it may be possible to achieve the event segmentation automatically using the PSMs’ kinematic data. The task board was static, such that the beginning and end of each tower interaction would occur near fixed coordinates. Additionally, the interaction with each tower involved only one hand, while the movements made between towers, which consisted of transferring the ring between the hands, involved both hands. Lastly, the order of the four towers remained constant. Hence, the beginning of the interaction with a tower could potentially be identified by correlating the arrival of the relevant PSM to the start area of the tower with near-zero velocity in the other PSM. Similarly, the end of the interaction with a tower would involve arriving at the end area of the tower, correlated with the other PSM beginning to move. These suggestions would increase the computational cost and would likely affect the performance to an extent; however, they offer potential avenues for fully automating detection using our algorithm.
This algorithm could be developed further to generalize to other training tasks. Although this would require adapting certain aspects of it, such as the threshold values, we posit that the general premise may still be applicable for detecting collision errors in various tasks. It may be possible to use different methods, including thresholding [doignon2004detection], YOLO [ragab2024comprehensive], and additional deep learning methods [kamtam2025deep] to locate the objects of interest. The movements of these objects can then be computed at the frame level using optical flow. Setting an appropriate threshold in the frequency domain can potentially provide a robust way of differentiating between occlusions and natural movements of anatomical features, and movements of the object of interest caused by collisions. Hence, although our algorithm includes task-specific aspects and thresholds, we hypothesize that it could potentially be expanded to other tasks and environments. Future work is needed to test this hypothesis.
V Conclusions
We studied the progression of surgical residents’ collision errors in the ring tower transfer task over a six-month period. To accomplish this, we developed an algorithm for the automatic detection of the collision errors, enabling faster and more consistent annotations. We additionally computed the task completion time and found that surgical residents learned to perform the ring tower transfer task faster over the span of six months. Many residents also reduced their errors, reflecting learning and improvement, while for several, shorter task times came at the expense of increased errors. Beyond enabling us to study the characteristics of the residents’ learning process, this framework provides a labeled dataset for further analysis.